a first order reaction has a rate constant k= 1.9*10^-4 s^-1. what is the half life for this reaction
1
Expert's answer
2018-01-26T14:36:08-0500
K = 1/t ln C0/C – formula for the first order reaction. For the half time calculations we use:
K = ln2/t1/2
t1/2 = ln2 / k = ln 2 / 1.9 ∙ 10-4 s-1 = 3648 s